OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

@brianna0214 slope point form is y-y1 = m(x-x1) where m = slope of the line and (x1,y1) points through which line line is passing hope u can do part1 now?

OpenStudy (loser66):

hehehe,,, if you want me the person shoot first, just speak out. no need to say that, guy Ok, I shoot

OpenStudy (loser66):

ok, slope =4 it means m =4 . so the slope form is y = 4x +b now, plug the point (3,-6) into the form to find b . we got -6 = 4 * 3 +b ---> b = -6 - 12 = -18 Now, plug b back to slope form to get y = 4x -18 Me done now , your turn for part b)